# Approximate Nearest Neighbor Search with Window Filters

### v1

@@ Appendix B ChatGPT Queries for RedCaps Query Generation
Appendix C DiskANN Primer
DiskANN (Jayaram Subramanya et al., 2019) is an approximate nearest neighbor algorithm that builds a graph on the input dataset $D$.
@@ Appendix C DiskANN Primer
To construct the “fast-preprocessing” variant, which is the variant used in practice, we start with an empty graph. We then do a beam search query for the nearest neighbor for all points $x\in D$, twice, building up the graph as we go. For each search, we record all points traversed in the beam search, along with the nearest neighbor if it wasn’t found, and then add these as neighbors to $x$. We then prune this list using the same heuristic we use for the “slow-preprocessing” variant, and also enforce with a hard cutoff that there are at maximum degree number of neighbors.
A beam search of size $B$ is a generalization of a greedy search. Given a query $x$, we start at a start node $s$ and “explore” $s$ by adding neighbors of $s$ to a queue. This queue consists of only the closest $B$ points to $x$ we have seen so far, explored or unexplored. We continually explore the closest unexplored node from the queue to $x$ until all nodes in the queue are explored. By increasing $B$, the beam search explores more points and is more likely to find a better nearest neighbor of $x$. We do beam searches for “fast preprorocessing” index construction, as described above, and for approximate nearest neighbor queries.
@@ D.2 Pareto Frontiers
See Figure 6, Figure 7, and Figure 8 for full Pareto frontier plots for a representative sample of filter widths on all datasets not included in thr main text.
@@
D.3 Branching Factor
Figure 9: Plots of index size and build time for varying branching factors on SIFT. The indices were built using 96 threads. Smaller values are better.
Figure 10: Pareto curves of recall vs. throughput on SIFT for varying window sizes and branching factors. The experiment was run using 16 threads. Up and to the right is better.
See Figure 9 for the index size and build times for different values of $\beta$ on SIFT. See Figure 10 for comparisons of varying $\beta$ on SIFT across all filter fractions. See Table 6 for speedups of our best method over the best baseline across recall levels $0.8$, $0.9$, $0.99$, and $0.995$.
@@ D.3 Branching Factor
Table 6: Speedup of our best method over the best baseline, restricted to hyperparameter settings that yield the recall in parenthesis in the dataset column. N/A means that none of our methods achieved that recall (on Redcaps this is due to poor DiskANN graph quality).

### v2

@@ Appendix B ChatGPT Queries for RedCaps Query Generation
Appendix C Vamana Primer
Vamana (Jayaram Subramanya et al., 2019) is an approximate nearest neighbor algorithm that builds a graph on the input dataset $D$. The entire system from (Jayaram Subramanya et al., 2019), including checkpointing and error recovery, is called DiskANN; Vamana is solely the in-memory ANNS component.
@@ Appendix C Vamana Primer
To construct the “fast-preprocessing” variant, which is the variant used in practice, we start with an empty graph. We then do a beam search query for the nearest neighbor for all points $x\in D$, twice, building up the graph as we go. For each search, we record all points traversed in the beam search, along with the nearest neighbor if it was not found, and then add these as neighbors to $x$. We then prune this list using the same heuristic we use for the “slow-preprocessing” variant, and also enforce with a hard cutoff that there are at maximum degree number of neighbors.
A beam search of size $B$ is a generalization of a greedy search. Given a query $x$, we start at a start node $s$ and “explore” $s$ by adding neighbors of $s$ to a queue. This queue consists of the closest $B$ points to $x$ we have seen so far, explored or unexplored. We continually explore the closest unexplored node from the queue to $x$ until all nodes in the queue are explored. By increasing $B$, the beam search explores more points and is more likely to find a better nearest neighbor of $x$. We do beam searches for “fast preprorocessing” index construction and for approximate nearest neighbor queries.
@@ D.2 Pareto Frontiers
See Figure 6, Figure 7, and Figure 8 for full Pareto frontier plots for a representative sample of filter widths on all datasets not included in the main text.
@@
D.3 Comments on Memory and Performance
We expect our method to scale well to larger datasets because our theory predicts only a $\log_{\beta}N$ factor increase in memory cost over a single Vamana index constructed on $D$. Our implementation also addresses memory size further by not constructing tree nodes that represent subsets smaller than $1000$ points and only storing the dataset $D$ once. Figure 9 shows the memory and construction time of a $\beta$-WST tree constructed for SIFT as we increase $\beta$. For large $\beta$, a $\beta$-WST tree has an only slightly larger memory footprint (about 2X) than a single ANN index, and for the theoretical and implementation reasons above we expect this trend to hold for larger datasets.
Finally, we run additional performance experiments. See Figure 10 for comparisons of varying $\beta$ on SIFT across all filter fractions, and see Table 6 for speedups of our best method over the best baseline across recall levels $0.8$, $0.9$, $0.99$, and $0.995$.
Figure 9: Plots of index size and build time for varying branching factors $\beta$ for $\mathsf{VamanaWST}$ on SIFT. The indices were built using 96 threads. Smaller values are better.
Figure 10: Pareto curves of recall vs. throughput on SIFT for varying window sizes and branching factors $\beta$ for $\mathsf{VamanaWST}$. The experiment was run using 16 threads. Up and to the right is better.
Table 6: Speedup of our best method over the best baseline, restricted to hyper-parameter settings that yield the recall in parenthesis in the dataset column. N/A means that none of our methods achieved that recall (on Redcaps this is due to poor Vamana graph quality).